# Quantum Teleportation Coexisting with Conventional Classical Communications in Optical Fiber

### v2

@@ 3.1 Noise rates, entanglement distribution, and Hong-Ou-Mandel interference
We now evaluate the performance of our quantum system. First, we characterize how much noise is introduced into the detectors $D_{1}$ and $D_{2}$ as a function of the classical power $P_{\text{cl}}$ that is multiplexed into the 30.2-km fiber. Fig. 2(a) shows the single photon count rates as we increase $P_{\text{cl}}$, where the SNSPD dark count rates have been subtracted. The extra noise scales linearly with power as 79.0 counts/s/mW and 97.9 counts/s/mW in $D_{1}$ and $D_{2}$, respectively, which differ due to slightly different detection system efficiencies. These results demonstrate that SpRS noise is significantly suppressed, as even mW power levels introduce extraneous noise on the order of the SNSPD dark count level. Since the 400-Gbps system can operate at $P_{\text{cl}}>0.5$ mW, this means that the SNSPDs hardly notice any rise in background counts when the system is operating at the minimum necessary classical power.
@@
Fig. 2(b) shows how Bob’s entanglement source’s non-local two-photon interference fringe visibility in the vertical and anti-diagonal bases is impacted after distributing one photon over 15.0 km to Charlie as we increase the classical power launched into the full 30.2-km link. The visibility is defined as $V_{\rm ent}=(R_{\rm max}-R_{\rm min})/(R_{\rm max}+R_{\rm min})$, where $R_{\rm max(\rm min)}$ are the maximum (minimum) two-fold coincidence counts of the interference fringe. Fig. 2(b) shows no noticeable degradation of $V_{\rm ent}$ up to $P_{\text{cl}}=$ 74 mW, in which we measure visibilities of $V_{V}=97.5\pm 0.1\%$ and $V_{A}=95.3\pm 0.2\%$. All error bars in this paper are determined via the Monte-Carlo method assuming Poisson photon counting statistics. These results indicate that entanglement distribution is well preserved at high classical powers, and that visibility is mainly limited from unity due to multi-pair emission during SPDC. These values are also well above the $1/\sqrt{2}$ bound for demonstrating the non-local nature of quantum entanglement [47]. Thus, this state is more than sufficient to use as a resource for quantum teleportation.
Next, we characterize HOM interference between heralded photons from Alice and Bob’s sources as we increase $P_{\text{cl}}$. To ensure polarization indistinguishably, we align the PBSs in the BSM setup to instead project onto $|H\rangle_{D_{1}}|H\rangle_{D_{2}}$ and prepare Alice and Bob’s heralded photons in the state $|H\rangle$. We then record four-fold coincidences as we vary the relative time-of-arrival of each photon using Alice’s variable optical delay. The quality of the interference is determined by the visibility of the fringe $V_{\rm HOM}=(R_{\rm max}-R_{\rm min})/R_{\rm max}$, where $R_{\rm max(min)}$ are the maximum (minimum) four-fold coincidence rates of the HOM interference fringe. The results are shown in Fig. 2(c). We measure a visibility of $V_{\rm HOM}=(R_{\rm max}-R_{\rm min})/R_{\rm max}=82.9\pm 4.5\%$ without the classical source, which is comparable to the $V_{\rm HOM}=83.1\pm 1.4\%$ visibility we obtain without distribution over the 30.2-km fiber. We measure $V_{\rm HOM}=80.3\pm 3.8\%$ with $P_{\text{cl}}$ set to 74 mW, showing interference is not significantly degraded by high-power coexisting C-band light. Each visibility is well above the classical bound of 50% [48], indicating our photons are sufficiently indistinguishable for performing quantum teleportation.
Figure 3: Four-fold coincidence counts as we scan Bob’s polarization analyzer setting along the Bloch sphere for Alice transmitting $|{D}\rangle$ (blue) and $|A\rangle$ (red) when the launch power of the 400-Gbps classical signal into the 30.2-km fiber is set to 74 mW.
Figure 4: Real and imaginary components of the density matrices obtained by single-qubit quantum state tomography on Bob’s target photon, conditioned on a successful BSM, whilst the launch power of the 400-Gbps C-band signal into the 30.2-km fiber is set to 74 mW. (a-d) show the results for Alice transmitting $|\psi\rangle_{A}=|H\rangle$, $|V\rangle$, $|D\rangle$, and $|A\rangle$, respectively.
@@ 3.2 Quantum teleportation coexisting with high-rate classical communications
Looking at the performance of entanglement distribution and HOM interference, we find that the key operations underlying quantum teleportation are well maintained at high C-band power levels. To demonstrate that this translates to the ability to perform noise-resilient teleportation, we evaluate the teleportation of various qubits while simultaneously transmitting 74 mW of C-band classical power.
@@
Fig. 3 shows four-fold coincidence counts as we scan Bob’s polarization analyzer setting along the Bloch sphere for Alice transmitting $|{D}\rangle=\frac{1}{\sqrt{2}}(|{H}\rangle+|{V}\rangle)$ and $|{A}\rangle=\frac{1}{\sqrt{2}}(|{H}\rangle-|{V}\rangle)$. The coincidences exhibit sinusoidal fringes that coincide with Alice’s prepared qubit. We obtain fringe visibilities of $V_{D}=81.3\pm 5.4\%$ and $V_{A}=74.7\pm 4.7\%$, both values being above the $V>1/3$ classical limit [49].
@@
Next, we perform maximum likelihood quantum state tomography [50] on Bob’s target photon to reconstruct the density matrix $\rho_{B}$ of his quantum state conditioned on a BSM detection. Fig. 4 shows $\rho_{B}$ for Alice transmitting the states $|H\rangle$, $|V\rangle$, $|D\rangle$, and $|A\rangle$. The fidelity of $\rho_{B}$ to Alice’s ideal qubit $|\psi\rangle_{A}$ is determined by calculating $F_{\psi}=\langle\psi_{A}|\rho_{B}|\psi_{A}\rangle$. State transfer in the H/V basis does not require HOM interference due to the projection onto $|H\rangle_{D_{1}}|V\rangle_{D_{2}}$ in the Bell state analyzer [51], where raw fidelity is mainly limited from unity due to multi-photon pair degradation. In this basis, we measure $F_{H}=97.5\pm 1.2\%$ and $F_{V}=95.8\pm 2.5\%$ for $|{\psi}\rangle_{A}=|{H}\rangle$ and $|{V}\rangle$, respectively. However, transferring coherent superpositions of $|H\rangle$ and $|V\rangle$ requires interference and thus exhibit more decoherence due to an imperfect spectro-temporal indistinguishability between Alice and Bob’s photons. For Alice transmitting $|{\psi}\rangle_{A}=|{D}\rangle$ and $|{A}\rangle$, we obtain fidelities of $F_{D}=87.5\pm 3.9\%$ and $F_{A}=85.5\pm 3.7\%$, respectively. Knowing that SpRS noise rates are independent of Alice’s prepared qubit, assuming symmetry of the equatorial states we calculate the average fidelity for an arbitrary qubit via $F_{\rm avg}=(F_{H}+F_{V}+2(F_{D}+F_{A}))/6$ in which we obtain $F_{\rm avg}=89.9\pm 3.1\%$. Furthermore, we measure $F_{\rm avg}=90.4\pm 2.7\%$ without sending the photons over the 30.2-km link, demonstrating that there is minimal degradation of fidelity from distribution over the fiber link nor the inclusion of the classical light. All of the measured fidelities are shown in Fig. 5. Since all are well above the classical bound of $F=2/3$ [49], we have clearly demonstrated that non-classical teleportation can be performed alongside high-rate conventional classical communications.
Figure 5: Fidelity of the reconstructed density matrix of Bob’s state to Alice’s initially prepared state for four different qubits as well as the average fidelity when the power of the 400-Gbps classical signal is set to 74 mW. The horizontal line shows the $F>2/3$ limit for demonstrating non-classical teleportation.
@@ 4 Discussion and Outlook
Since teleportation was conducted alongside much higher powers than necessary for the 400-Gbps classical system, this indicates that substantially higher data rates could have propagated through the fiber. The classical power used was 148 times greater than necessary for a single 400-Gbps channel, and 870 times greater than if the receiver were located directly at the end of the 30.2-km fiber. Thus, a multi-channel WDM classical system could easily have carried many terabits/s aggregate data rates without having a significant impact on quantum fidelity. The demonstrated powers are also on the order of the highest achieved in state-of-the-art single [10, 11, 12] and entangled [17] photon coexistence studies. This indicates the potential for teleportation-based quantum networking systems, such as quantum repeaters, operating alongside classical backbone networks [10].
Since our system uses two different photon pair sources, multiple nodes, and four-fold coincidence detection, our results indicate the potential for more complex systems operating alongside classical networks. For example, simple modifications to Alice’s source could allow entangled photon pair generation. A BSM at Charlie would then result in entanglement swapping [52], establishing entanglement between initially independent photons at Alice and Bob’s nodes. Due to our system’s high tolerance to classical power, it is reasonable to conclude that a similarly designed entanglement swapping system could operate alongside high-rate coexisting classical communications. Further, our multi-node configuration with the BSM performed near the midpoint of a longer fiber link is a key element for multi-node networking with a central quantum processor [34] and to implement quantum repeaters [29]. We note that these applications will also require quantum memories to store incident photons, the degree to which SpRS noise will impact various platforms and applications will require future investigation.
The state teleportation fidelity we obtained is comparable to recent experiments conducted over quantum-dedicated fiber networks [34, 33, 37], but could be further improved via standard methods. We showed that opposed to SpRS noise from coexisting classical light, our fidelity was mainly limited by multi-photon pair emission during SPDC and imperfect photon indistinguishably during the BSM. Raw fidelity could thus be improved by reducing the photon pair generation probability [46] and increasing indistinguishability by decreasing the ratio between the source’s pump spectral bandwidth to the photon filter bandwidths [42]. When considering SpRS noise, narrower filters would simultaneously increase indistinguishability and reject more SpRS noise. Altogether, employing these approaches could allow our teleportation fidelity to approach more closely to unity and have an even higher tolerance to classical power.
@@
Finally, classical signals are often used for quantum networking purposes. Some applications include communicating measurement results between nodes (eg., the BSM), synchronizing distant pulsed sources and detectors to a common clock, and channel monitoring. Similarly, quantum-classical coexistence can enhance functionality without sacrificing potentially limited fiber resources [16, 53, 22, 21]. Since these signals can typically operate well below the powers demonstrated here (sub-mW levels), our results show that including these signals in our system would have negligible impact if allocated to the C- or L-bands.
@@
In conclusion, we have demonstrated a three-node quantum state teleportation system operating in a 30.2-km fiber link that simultaneously carries 400-Gbps conventional classical communications. By employing various methods to suppress the impact of SpRS noise photons, teleportation fidelity was well maintained with elevated classical powers capable of transmitting many terabits/s aggregate data rates. Since we have clearly shown that appropriately designed teleportation systems can operate alongside modern high-rate data traffic, this work demonstrates a significant step towards ensuring future quantum networking can be realized anywhere in the existing fiber infrastructure.
@@
Acknowledgments The authors would like to acknowledge the support of StarLight International/National Communications Exchange Facility and Ciena Corporation. We further thank the members of the Advanced Quantum Networks for Scientific Discovery (AQNET-SD) project, which is a collaboration between Northwestern University, Caltech, Fermilab National Laboratory, Argonne National Laboratory, the Jet Propulsion Laboratory, and University of Illinois Urbana-Champaign towards advancing quantum network technology and applications.

### v3

@@ 3.1 Noise rates, entanglement distribution, and Hong-Ou-Mandel interference
We now evaluate the performance of our quantum system. First, we characterize how much noise is introduced into the detectors $D_{1}$ and $D_{2}$ as a function of the classical power $P_{\text{cl}}$ that is multiplexed into the 30.2-km fiber. Fig. 2(a) shows the single photon count rates as we increase $P_{\text{cl}}$, where the SNSPD dark count rates have been subtracted. The noise rates scale linearly as 79.0 counts/s/mW and 97.9 counts/s/mW in $D_{1}$ and $D_{2}$, respectively, the difference is due to slightly unequal efficiencies after the 50:50 coupler. Since a single 400-Gbps channel can operate with as low as $P_{\text{cl}}=0.5$ mW, the system can operate with nearly no noticeable noise increase above the SNSPD dark count rate. Fig. 2(a) also includes an axis showing the ratio $M$ of $P_{\text{cl}}$ to the minimum single-channel power, which can be used to estimate the maximum aggregate data rates the $P_{\text{cl}}$ could support in a multi-channel system via $M\times 400$ Gpbs.
@@
Fig. 2(b) shows how Bob’s entanglement source’s non-local two-photon interference fringe visibility in the vertical and anti-diagonal bases is impacted after distributing one photon over 15.0 km to Charlie as we increase the C-band signal’s power launched into the full 30.2-km link. The visibility is defined as $V_{\rm ent}=(R_{\rm max}-R_{\rm min})/(R_{\rm max}+R_{\rm min})$, where $R_{\rm max(\rm min)}$ are the maximum (minimum) two-fold coincidence counts of the interference fringe. Fig. 2(b) shows no noticeable degradation of $V_{\rm ent}$ up to $P_{\text{cl}}=$ 74 mW, in which we measure $V_{\rm ent,V}=97.5\pm 0.1\%$ and $V_{\rm ent,A}=95.3\pm 0.2\%$. All error bars in this paper are determined via the Monte-Carlo method assuming Poisson photon counting statistics. These results indicate that entanglement distribution is well preserved and that visibility is mainly limited by multi-pair emission during SPDC. These values are also well above the $1/\sqrt{2}$ bound for demonstrating the non-local nature of quantum entanglement [50] and more than sufficient to use for quantum teleportation.
Next, we characterize HOM interference between heralded photons from Alice and Bob’s sources as we increase $P_{\text{cl}}$. To ensure polarization indistinguishably, we align the PBSs in the BSM setup to instead project onto $|H\rangle_{D_{1}}|H\rangle_{D_{2}}$ and prepare Alice and Bob’s heralded photons in the state $|H\rangle$. We then record four-fold coincidences as we vary the relative time-of-arrival of each photon using Alice’s variable optical delay. The quality of the interference is determined by the visibility of the fringe $V_{\rm HOM}=(R_{\rm max}-R_{\rm min})/R_{\rm max}$, where $R_{\rm max(min)}$ are the maximum (minimum) four-fold coincidence rates of the HOM interference fringe. The results are shown in Fig. 2(c). We measure a visibility of $V_{\rm HOM}=82.9\pm 4.5\%$ without the classical source, which is comparable to the $V_{\rm HOM}=83.1\pm 1.4\%$ visibility we obtain without distribution over the 30.2-km fiber (see supp. materials). We measure $V_{\rm HOM}=80.3\pm 3.8\%$ with $P_{\text{cl}}$ set to 74 mW, showing interference is not significantly degraded by high-power coexisting C-band light. Each visibility is well above the classical bound of 50% [51], indicating our ability to demonstrate non-classical interference across the range of the tested coexisting C-band powers.
@@ 3.2 Quantum teleportation coexisting with high-rate classical communications
Figure 3: Four-fold coincidence counts as we scan Bob’s polarization analyzer setting around the Bloch sphere for Alice transmitting $|{D}\rangle$ (blue) and $|A\rangle$ (red) when the launch power of the 400-Gbps C-band classical signal into the 30.2-km fiber is set to 74 mW.
@@
Looking at the performance of entanglement distribution and HOM interference, we find that the key operations underlying quantum teleportation are well maintained at high C-band power levels. To demonstrate that this translates to the ability to perform teleportation, we evaluate the teleportation of various qubits while simultaneously transmitting 74 mW of C-band classical power.
@@
Fig. 3 shows four-fold coincidence counts as we scan Bob’s polarization analyzer setting around the Bloch sphere for Alice transmitting $|{D}\rangle=\frac{1}{\sqrt{2}}(|{H}\rangle+|{V}\rangle)$ and $|{A}\rangle=\frac{1}{\sqrt{2}}(|{H}\rangle-|{V}\rangle)$. The coincidences exhibit sinusoidal fringes that coincide with Alice’s prepared qubit. We obtain fringe visibilities of $V_{D}=81.3\pm 5.4\%$ and $V_{A}=74.7\pm 4.7\%$, both values being above the $V=1/3$ classical limit [52].
@@
Next, we perform maximum likelihood quantum state tomography [53] on Bob’s target photon to reconstruct the density matrix $\rho_{B}$ of his quantum state conditioned on a BSM detection. Fig. 4 shows $\rho_{B}$ for Alice transmitting the states $|H\rangle$, $|V\rangle$, $|D\rangle$, and $|A\rangle$. The fidelity of $\rho_{B}$ to Alice’s ideal qubit $|\psi\rangle_{A}$ is determined by calculating $F_{\psi}=\langle\psi_{A}|\rho_{B}|\psi_{A}\rangle$. State transfer in the H/V basis does not require HOM interference due to the projection onto $|H\rangle_{D_{1}}|V\rangle_{D_{2}}$ in the Bell state analyzer [54], where raw fidelity is mainly limited from unity due to multi-photon pair degradation. In this basis, we measure $F_{H}=97.5\pm 1.2\%$ and $F_{V}=95.8\pm 2.5\%$ for $|{\psi}\rangle_{A}=|{H}\rangle$ and $|{V}\rangle$, respectively. However, transferring coherent superpositions of $|H\rangle$ and $|V\rangle$ requires interference and thus exhibit more decoherence due to an imperfect spectro-temporal indistinguishability between Alice and Bob’s photons. For Alice transmitting $|{\psi}\rangle_{A}=|{D}\rangle$ and $|{A}\rangle$, we obtain fidelities of $F_{D}=87.5\pm 3.9\%$ and $F_{A}=85.5\pm 3.7\%$, respectively. Knowing that SpRS noise rates are independent of Alice’s prepared qubit, assuming symmetry of the equatorial states we calculate the average fidelity for an arbitrary qubit via $F_{\rm avg}=(F_{H}+F_{V}+2(F_{D}+F_{A}))/6$ in which we obtain $F_{\rm avg}=89.9\pm 3.1\%$. Furthermore, we measure $F_{\rm avg}=90.8\pm 0.8\%$ without sending the photons over the 30.2-km link (see supp. materials), demonstrating that there is minimal degradation of fidelity from distribution over the fiber nor the inclusion of the 400-Gbps signal. The state transfer fidelity of each qubit is shown in Fig. 4(e). All are well above the $F=2/3$ limit for classical-based methods [52], demonstrating that non-classical teleportation can be achieved alongside high-rate conventional communications.
Figure 4: Real and imaginary components of the density matrices obtained by single-qubit quantum state tomography on Bob’s target photon, conditioned on a successful BSM, whilst the launch power of the 400-Gbps C-band signal into the 30.2-km fiber is set to 74 mW. (a-d) show the results for Alice transmitting $|\psi\rangle_{A}=|H\rangle$, $|V\rangle$, $|D\rangle$, and $|A\rangle$, respectively. (e) The fidelity of Bob’s density matrix to Alice’s ideally prepared qubit and the calculated average fidelity. The dotted line shows the $F=2/3$ limit for the highest fidelity achievable using only classical physics-based methods [52].
@@ 4 Discussion and Outlook
In this paper, teleportation was performed alongside $M\approx$150 times higher classical powers than necessary for error-free 400-Gbps communications over the 78-km fiber and $M\approx$870 times greater for a 30.2-km link. Thus, WDM communications could have easily reached many Tbps aggregate data rates without significant impact on teleportation fidelity. The classical power was also on the order of the highest achieved in state-of-the-art studies using direct transmission of quantum states [25, 10, 11, 12, 17] and indicates the potential for teleportation-based applications operating in the classical backbone fiber infrastructure [10].
Our results also indicate the potential for integrating even more advanced quantum operations. For example, simple modifications to Alice’s photon pair source would generate entangled photon pairs and result in entanglement swapping [55], entangling initially independent photons at Alice and Bob’s nodes after the BSM. Based on this similarity, it is reasonable to conclude that swapping alongside high-rate classical data is achievable. Further, midpoint BSMs are important for applications such as networking with a central quantum processor [34], quantum repeaters [29], amongst others. Although these will require more capabilities beyond those demonstrated here (eg., quantum memories, purification, etc.), our results are promising and could be useful for the ongoing development of this technology.
Although O-band teleportation networks show promise for high-fidelity integration, C-band quantum networking is usually preferred in quantum-dedicated fiber due to the lowest loss in silica fibers. Unfortunately, C-band channels suffer from orders of magnitude higher SpRS generation rates from C-band classical light [4, 17]. CV quantum encoding can help address this issue [25] but is much more fragile to loss compared to discrete-variables which limits its reach. One intriguing possibility is to use high-dimensional entanglement [56], but has yet to be fully investigated in high-power regimes. Using a theoretical model [48], we predict our teleportation fidelity would be significantly impacted with mW-level C-band classical power if we used C-band quantum signals (see supp. materials), but coincidence rates would roughly triple. This has relevance for research towards long-distance quantum backbone networks since conventional backbone fibers can have $\sim 100$ mW aggregate power. Further, WDM bandwidth may become limited as some conventional networks push towards fully populating the C- and L-bands [57], which also increases the difficulty of avoiding channel cross-talk and optical amplifier noise. One route to alleviate lower O-band quantum rates is to use wavelength multiplexing, which has been shown successful in allowing multi-channel entangled photons populating most of the O-band co-propagating with high-power C- and L-band WDM channels [17]. In any case, future research into novel methods to protect quantum fidelity from strong background noise would be invaluable.
Some teleportation-based applications require storing photons in quantum memories (QMs), for example to store target photons long enough to receive a signal encoding the BSM result and apply the corresponding unitary transformation to recover Alice’s exact state before usage. In our experiment we performed a posteriori teleportation [58] since Bob’s photon is detected before the BSM occurs. Interestingly, quantum mechanics predicts the same correlations irrespective of measurement order. However, attempting to store photons over a long duration can increase photon loss which may complicate the dependence on measurement order in real-world settings having appreciable background noise. If Bob’s target channel is completely noiseless, our model predicts that adding pure loss to his target photon’s channel does not further degrade the fidelity set by the noise level in the BSM detectors. However, this no longer holds if there is some noise count probability in Bob’s detector. In our case we predict our low dark count SNSPDs could tolerate roughly 60 dB loss in the target channel, which is more than manageable using spooled fiber to delay measurement. However, other approaches such as those using quantum frequency conversion to interface telecom wavelength photons with particular QMs can have notable noise-levels [37] that place more stringent requirements on Bob’s channel efficiency (see supp. materials).
It is interesting to consider whether teleporting quantum information has any unique benefits in quantum-classical networking compared to direct transmission. Interestingly, we predict that our system can achieve higher coexisting classical power compared to if Alice had directly sent her qubit to Bob over the fiber using a WCS or a heralded single-photon (see supp. materials). This could be due to the stronger correlations in detection across the four photons in our system, and also the segmenting of the longer fiber with a midpoint BSM [28]. Future studies will compare with other teleportation designs such as those using a WCS source at Alice’s node [33, 36, 39] instead of our heralding approach. Looking beyond state teleportation, it would be interesting to investigate the impact of coexisting classical channels on concatenated entanglement swapping or quantum repeaters.
@@
In regards to real-world deployment, classical light-based signals are often used in quantum networking for synchronizing distant pulsed quantum sources and detectors to a common clock, communicating measurement results (eg., BSM results), and monitoring environmental disturbances. Similarly, multiplexing these signals can enhance functionality without sacrificing potentially limited fiber resources [16, 59, 22, 21]. Since these can typically operate at sub-mW powers, our results show that including these in our system would have negligible impact if allocated to the C- or L-bands.
@@
In conclusion, we have demonstrated quantum state teleportation over a 30.2-km fiber that is populated with high-power 400-Gbps conventional data traffic. By employing various methods to suppress SpRS noise, teleportation fidelity was well maintained alongside elevated classical powers capable of transmitting many Tbps aggregate data rates. We further investigated multiple key questions for optimizing teleportation-based applications and identified challenges facing the deployment of future quantum networks. Altogether, this work demonstrates a significant step towards ensuring complex multi-photon/multi-node quantum network applications can be realized anywhere in the existing fiber infrastructure.
@@
Acknowledgments The authors would like to acknowledge the support of StarLight International/National Communications Exchange Facility and Ciena Corporation. We further thank the members of the Advanced Quantum Networks for Scientific Discovery (AQNET-SD) project, which is a collaboration between Northwestern University, Caltech, Fermilab, Argonne National Laboratory, Jet Propulsion Laboratory, and University of Illinois Urbana-Champaign. We also thank Gina Talcott for help in the editing process.